OpenStudy (anonymous):

anyone...Pls help A commuter travels to work, driving a distance d at speed r. At the end of the day, she returns home and drives the same distance 4 mph faster, at a speed r+4. Her average speed for the round trip can be represented by the expression 2d/(d/r)+(d/r+4)

OpenStudy (dumbcow):

avg speed = total distance/total time d = r*t => t=d/r total distance = 2d total time = d/r + d/r+4 = (d(r+4) + dr)/r(r+4) = (2dr+4d)/r(r+4) avg speed = 2d*(r(r+4)/2d(r+2) = r(r+4)/(r+2) wait what are we trying to solve for...

OpenStudy (anonymous):

it also says simplify this expression and then gives me 4 choices as follows

OpenStudy (anonymous):

r(r+4)/(r+2) or r(r+2)/(r+4) or dr(r+4)/(dr+2) or 1/ (r+2)

OpenStudy (dumbcow):

ok then look at what i did above A)

OpenStudy (anonymous):

yep...thanks you r awesome